OpenStudy (anonymous):

what is the answer celeste s air conditioner uses 2160w of power as a current of 9.0 A passes through it. a) what is the voltage drop when the air conditioner is running? b) How does this compare to the usual house voltage? c) What would happen if Celeste tried connecting her air conditioner to a usual 120-V line?

OpenStudy (anonymous):

P=V*I V=P/I=2160/9 V=240V household voltage is 120V if V_1=V/2, for the power to be the same the current must be...(P=V*I)